OpenStudy (anonymous):

If circle P is tangent to both the x-axis and the y-axis and the coordinates of the center are (r,r) what is the length of the chord whose endpoints are the points of tangency? and what is the equation of the line containing the chord?

OpenStudy (anonymous):

a. \[r \sqrt{2}\] y=x-1

OpenStudy (anonymous):

not sure with eqn

OpenStudy (anonymous):

sure if the circle is in 4th quadrant

OpenStudy (anonymous):

Assuming first quadrant, r sqrt 2 and y= -x + r